OpenStudy (anonymous):

need help with diff. eq. h/w y'' +4y = sin^2(x) need help with diff. eq. h/w y'' +4y = sin^2(x) @Mathematics

OpenStudy (anonymous):

I found the roots and have the complementary equations y=c1cos(2x)+c2sin(2x) where c1 and c2 are constants

OpenStudy (jamesj):

So there are two steps here ... and you've done the first step.

OpenStudy (anonymous):

then I took the derivative and double derivative of y

OpenStudy (jamesj):

Now for the particular solution, I recommend using the method of undetermined coefficients. Use a trial solution of yp = A sin^2 x + B sin x cos x + C cos^2 x

OpenStudy (anonymous):

Okay that's what i want to do but i am supposed to use a different method. where I assume part of the first derivative of y is equal to 0 and then plug in y^I and y^II into the original equal so I end up with 2 equations with two unknowns. I think I am supposed to assume the constants are a function of x

OpenStudy (anonymous):

I don't know what the method is called though

OpenStudy (jamesj):

You're meant to turn this into a first order system? That's also a very good method.

OpenStudy (anonymous):

i found 1 equation just by plugging in y'' +4y = sin^2(x) and cancelling some terms out. but I am confused how to find the second one

OpenStudy (jamesj):

What you mean in that case is that not powers of y as you've written, but define two new variables: y1 = y y2 = y' = y1' Then the equation is y'' + y = sin^2 x is equivalent to y2' + y1 = sin^2 x Hence the two equations of the first order system are y1' = 0 + y2 + 0 y2' = -y1 + 0 + sin^2 x

OpenStudy (anonymous):

not exactly the method that my teacher used in class was alil different

OpenStudy (anonymous):

you still solve for a particular and complementary equation and add then in the end

OpenStudy (jamesj):

Are you using perhaps the method of undetermined parameters? Using your two complimentary solutions, call them y1 and y2 (nothing to do with the y1 and y2 of what I just wrote) and finding a particular solution of the form: yp = u1(x) y1 + u2(x) y2 ?

OpenStudy (anonymous):

yes i think so that seem right

OpenStudy (jamesj):

Ok, so yp = u1 . sin 2x + u2 cos 2x Now differentiate twice and substitute into your equation. Once you've done that carefully, what do you get?

OpenStudy (anonymous):

I ened up with -2(u1)'sin(2x)+2(u2)'cos(2x)=sin^2x

OpenStudy (anonymous):

actually i might have u1 and u2 backwardds

OpenStudy (anonymous):

-2(u2)'sin(2x)+2(u1)'cos(2x)=sin^2x

OpenStudy (anonymous):

u1 and u2 are functions of x right?

OpenStudy (jamesj):

yes, most definitely.

OpenStudy (anonymous):

okay so if my first equation is right how do you find the second equation. I remember setting part of the first derivative of y equal to zero

OpenStudy (anonymous):

but that doesn't seem to work

OpenStudy (jamesj):

I don't agree with your equation above.

OpenStudy (jamesj):

Where are the second derivatives, for instance?

OpenStudy (anonymous):

the second derivative of y=u1cos2x+u2sin2x?

OpenStudy (jamesj):

Where are the second derivatives of the u1 and u2 ... are you sure they cancel out?

OpenStudy (jamesj):

yp = u1 . sin 2x + u2 cos 2x yp' = u1' sin 2x + u2' . cos 2x + 2 u1 . cos 2x - 2 u2 sin 2x yp'' = u1'' sin 2x + u2'' . cos 2x + 2 u1' . sin 2x - 2 u2' . cos 2x -4 u1 . sin 2x - 4 u2 . cos 2x + 2 u1' . cos 2x - 2 u2' . sin 2x Right? Hence yp'' + 4 yp = sin^2 x implies that ....

OpenStudy (jamesj):

stuff

OpenStudy (anonymous):

ohhh okay i see whats wrong when i did it i followed the example he did in class where he took the first derivative so in general you have y=c1(x)y(x)+c2(x)y2(x) y'=c1'(x)y(x)+c2'(x)y2(x)+c1(x)y'(x)+c2(x)y2'(x) then what he did was assume c1'(x)y(x)+c2'(x)y2(x)=0 and take the derivative of what is left in y'

OpenStudy (jamesj):

let me tell you the answer so you can work towards it. The particular solution is yp = 1/6 . cos 2x + 1/2

OpenStudy (anonymous):

okay that's helpful. But i am confused on the method I am supposed to use. My teacher had 2 equations c1'(x)y(x)+c2'(x)y2(x)=0 and whatever you you get when you plug into y'' +4y = sin^2(x

OpenStudy (anonymous):

where the second equation was in the form c1'(x)y'(x)+c2'(x)y2'(x)=f(x)

OpenStudy (jamesj):

I'm honestly not sure where that assumption c1'(x)y(x)+c2'(x)y2(x)=0 comes from right now; but use it for now as it will simplify your life a lot and see if it works. I'll think about why that might be true.

OpenStudy (jamesj):

Right.

OpenStudy (anonymous):

yeah i don't understand why that assumption works but from what I understood from class this method works on differential equation where you cant guess solutions from the method of undetermined coefficents

OpenStudy (jamesj):

Ok. I've got to run. But I'll think about this problem some more and comment when I come back.

OpenStudy (anonymous):

okay thanks its not due untill tomorrow at 11am so i might add some things if I figure anything out